2y^2+4x y dy/dx + xe^xy dy/dx + 2y dy/dx=-ye^xy
1
Expert's answer
2020-10-12T17:21:30-0400

"2y^2+4x y\frac{ dy}{dx} + xe^{xy} \frac{dy}{dx} + 2y\frac{ dy}{dx}=-ye^{xy}"


"(4x y + xe^{xy} + 2y)\frac{ dy}{dx}+2y^2+ye^{xy}=0"


"(2y^2+ye^{xy})dx+(4x y + xe^{xy} + 2y) dy=0"


"\frac{\partial}{\partial x}(4x y + xe^{xy} + 2y)=4y+e^{xy}+xye^{xy}"


"\frac{\partial}{\partial y}( 2y^2 + ye^{xy})=4y+e^{xy}+xye^{xy}"


Since "\frac{\partial}{\partial x}(4x y + xe^{xy} + 2y)=\frac{\partial}{\partial y}( 2y^2 + ye^{xy})", there exists a function "U=U(x,y)" such that "dU=(4x y + xe^{xy} + 2y) dy+(2y^2+ye^{xy})dx." Therefore, "\frac{\partial U}{\partial x}=2y^2 + ye^{xy}" and "\frac{\partial U}{\partial y}=4x y + xe^{xy} + 2y".


"U(x,y)=\int(2y^2 + ye^{xy})dx=2y^2x+e^{xy}+C(y)".


"4xy + xe^{xy}+2y=\frac{\partial U}{\partial y}=\frac{\partial }{\partial y}(2y^2x+e^{xy}+C(y))=4xy+xe^{xy}+C'(y)"


Therefore, "C'(y)=2y". So "C(y)=y^2+C".


We conclude that the general solution of the differential equation is the following:


"2y^2x+e^{xy}+y^2=C"